OpenStudy (anonymous):

A rectangle has width 40 in and length 24 in. Find the length of another rectangle of equal area, if the width is 30 in. A. 32 in B. 18 in C. 50 in

OpenStudy (anonymous):

40 * 24 = 30 * l 40 * 24 / 30 = l l = 32 A is the answer

hero (hero):

vyew?

OpenStudy (anonymous):

in an hour?

OpenStudy (anonymous):

pls :)

hero (hero):

I guess

OpenStudy (anonymous):

A

OpenStudy (anonymous):

u here hero??

hero (hero):

Lisa, you always seem to have something else going on when I am here so I end up helping other students.

OpenStudy (anonymous):

sorry :( my parents were talking to me about paying rent ...